OpenStudy (anonymous):

please help... The equation of line GH is y = 4x − 3. Write an equation of a line perpendicular to line GH in slope-intercept form that contains point (2, 3). a. y = 4x − 3 b.y = −4x − 5 c.11 over 4x − 3 d.y = negative 1 over 4x + 7 over 2

OpenStudy (triciaal):

perpendicular lines the product of the slopes is -1 given slope is 4 so the slope of the perpendicular is -1/4 2 points on the line are (x, y) and (2, 3) option 1 y-3 = -1/4(x - 2) y = -1/4(x) + 1/2 + 3 y = -1/4 (x) + 3.5 option 2 y = mx + b where m is the slopee using 2,3 3 = -1/4(2) + b so b = 3.5 for any x,y y = -1/4(x) + 3.5

OpenStudy (triciaal):

option d 3.5 = 7/2